Robert borrows $1000 for 8 months from a lender who charges a 16% discount rate. What is the equivalent simple interest rate he pays on the loan? Ans; 17.91%
Answer #1

The interest on 1000 = Borrowed amount * Discount rate * Time in year1000*16%*8/12 = 106.67
The discounted price = 1000 - 106.67 = 893.33

The equivalent simple = Interest/Discounted Price *(12/number of months) = 106.67/893.33 * (12/8) = 17.91%

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well